Using the elimination method, how would you solve the following
5x+5y=13
7x-2y=17

Answer by Alan3354(69443)

Using the elimination method, how would you solve the following
Using the elimination method, how would you solve the following
5x+5y=13
7x-2y=17
-----------
To eliminate either x or y, multiply the 2 eqns to get the same coefficients.
The y coeff's are smaller, so eliminate y's.
-----------
5x+5y=13 x2 --> 10x + 10y = 26
7x-2y=17 x5 --> 35x - 10y = 85
Add those 2, solve for x.
Then sub for x and find y.